# Examining Post-Training Quantization for Mixture-of-Experts: A Benchmark

### v1

@@ A.3 Further Discussion
Expert usage frequency.
As shown by Q1 in Section 4.2, expert usage frequency is a critical metric in the compression of MoE models, predicated on the insight that less frequently used experts are likely less crucial. We present further ablation discussion on the bits allocation in the expert-frequency-based methods.
Table 6: Ablation on the allocated bits for the selected top-$k$ experts based on frequency. We compare the allocation of {$4$, $8$} bits of the top-$k$ experts based on frequency, all other experts are quantized to $2$ bits.
@@
Figure 6: Performance of different quantization bits on DeepSeek-MoE-16B-base model.
@@
We summarize all experimental results and illustrate the relationship between bit width and average performance in Figure 6. Overall, we observe that as the bit width increases, the performance is improved. As highlighted by the red cross mark in the figure, achieving an average MoE bit width of $2.12$ results in a performance score of $61.11$, which marks a $5\%$ improvement over the model quantized to $2$ bits. This underscores the effectiveness of MoE blocks in settings with limited bit width.

### v2

@@ A.3. Further Discussion
Expert usage frequency.As shown by Q1 in Section 4.3, expert usage frequency is a critical metric in the compression of MoE models, predicated on the insight that less frequently used experts are likely less crucial. We present further discussion of ablation on the bits allocation in the expert-frequency-based methods.
Table 6. Ablation on the allocated bits for the selected top-$k$ experts based on frequency. We compare the allocation of {$4$, $8$} bits of the top-$k$ experts based on frequency, and all other experts are quantized to $2$ bits.
@@
Table 7. Baseline results of the $16$-bit (FP16161616), $4$-bit, and $2$-bit quantization.
Bits WinoGrande (%) COPA (%) OBQA (%) HellaSwag (%) PIQA (%) MMLU (%) Average (%)
Mixtral-8x7B
$16$ $76.48$ $93.00$ $47.00$ $83.98$ $82.37$ $70.35$ $75.33$
$4$ $74.98$ $92.00$ $46.20$ $81.65$ $80.85$ $67.65$ $73.89$
$2$ $49.33$ $63.00$ $25.40$ $28.18$ $52.99$ $24.29$ $40.53$
DeepSeek-MoE-16B-base
$16$ $70.40$ $91.00$ $44.20$ $77.35$ $78.72$ $44.77$ $67.74$
$4$ $71.35$ $87.00$ $43.20$ $76.39$ $78.51$ $44.22$ $66.78$
$2$ $53.28$ $76.00$ $30.20$ $45.33$ $66.54$ $25.28$ $49.44$
Combination of the weight outlier and expert usage frequency. We conducted additional experiments on the DeepSeek-MoE-16B-base model by integrating bit-width allocation based on layers with significant weight outliers with allocation based on expert usage frequency to explore the trade-off between them. Specifically, we aimed for a total average bit budget of $2.97$. We select portions of the model to be quantized to $4$ bits using a combination of the two heuristics, while quantizing all attention weights to $4$ bits and all other weights to $2$ bits. For selecting the $4$-bit weights, we introduce a hyper-parameter, $\alpha$ (00 < $\alpha$ < $1$), representing the proportion of weights chosen based on expert usage frequency, with the remainder selected based on weight outliers. We varies $\alpha$ to illustrate the trade-off between these methods, as detailed above. As shown in Table 8, the optimal combination of these two methods occurs when alpha is set to $0.1$. This means that $20\%$ of the $4$-bit MoE weights are selected based on expert usage frequency, while the remaining $80\%$ are chosen according to weight outliers.
Table 8. The combination of weight outlier and expert usage frequency, evaluated on the DeepSeek-MoE-16B-base model.
Bits $\alpha$ WinoGrande (%) COPA (%) OBQA (%) HellaSwag (%) PIQA (%) MMLU (%) Average (%)
$2.97$ $0.0$ $67.72$ $90.00$ $39.20$ $72.83$ $77.15$ $41.06$ $64.66$
$0.1$ $68.11$ $89.00$ $41.60$ $72.88$ $77.80$ $41.84$ $65.21$
$0.2$ $69.21$ $89.00$ $41.20$ $72.60$ $76.93$ $41.60$ $65.09$
$0.3$ $68.92$ $88.00$ $42.00$ $72.06$ $76.65$ $41.21$ $64.81$
$0.4$ $67.48$ $89.00$ $41.40$ $71.88$ $76.71$ $40.96$ $64.57$
$0.5$ $67.32$ $90.00$ $40.80$ $71.89$ $76.93$ $40.21$ $64.52$
$0.6$ $65.90$ $87.00$ $39.40$ $71.86$ $76.76$ $38.67$ $63.27$
$0.7$ $66.21$ $87.00$ $41.40$ $71.45$ $76.87$ $36.98$ $63.32$
$0.8$ $66.45$ $89.00$ $41.00$ $70.89$ $76.60$ $37.67$ $63.60$
$0.9$ $66.37$ $84.00$ $40.20$ $70.83$ $76.87$ $39.84$ $63.02$
$1.0$ $68.19$ $87.00$ $41.60$ $71.01$ $76.11$ $40.81$ $64.12$
Table 9. Baseline results of the $16$-bit (FP16161616), $4$-bit, and $2$-bit quantization.
Bits WinoGrande (%) COPA (%) OBQA (%) HellaSwag (%) PIQA (%) MMLU (%) Average (%)
Mixtral-8x7B
$16$ $76.48$ $93.00$ $47.00$ $83.98$ $82.37$ $70.35$ $75.33$
$4$ $74.98$ $92.00$ $46.20$ $81.65$ $80.85$ $67.65$ $73.89$
$2$ $49.33$ $63.00$ $25.40$ $28.18$ $52.99$ $24.29$ $40.53$
DeepSeek-MoE-16B-base
$16$ $70.40$ $91.00$ $44.20$ $77.35$ $78.72$ $44.77$ $67.74$
$4$ $71.35$ $87.00$ $43.20$ $76.39$ $78.51$ $44.22$ $66.78$
$2$ $53.28$ $76.00$ $30.20$ $45.33$ $66.54$ $25.28$ $49.44$
Baseline results of low-precision quantization. We provide the $16$-bit (FP16161616), $4$-bit, and $2$-bit baselines of both Mixtral-8x7B and DeepSeek-MoE-16B-base models in Table 9.